OpenStudy (anonymous):

Find the measure of each interior angle of a regular dodecagon. A. 30° B. 180° C. 150° D. 360°

OpenStudy (anonymous):

For any polygon, Interior angle=180 degrees-exterior angle. What is the exterior angle?

OpenStudy (anonymous):



OpenStudy (anonymous):

ohh 360

OpenStudy (anonymous):

no exterior angle = 360/#of sides

OpenStudy (anonymous):

30 degrees

OpenStudy (anonymous):

Good, you found the exterior angle. 360/12=30-----------360/(# of sides)=(exterior angle) and 180-30=150----------180-(exterior angle)=(interior angle) answer is 150

OpenStudy (anonymous):

now using exterior angle you can find interior angle.

OpenStudy (anonymous):

Interior Angle in a Polygon of side n = (n-2) × 180° / n A dodecagon has 12 sides, so n= 12. Interior Angle = (12-2) x 180 / 10 = 10x180 / 10 = ?

OpenStudy (anonymous):

thanks i think i got it now @rrrrr